# Modular Automatic Complexity Analysis of Recursive Integer Programs

### v1

@@ 1 Introduction
There exist numerous approaches to analyze complexity of programs automatically, e.g., [ben-amram2017MultiphaseLinearRankingFunctions, albert2019ResourceAnalysisDriven, sinn2017ComplexityResourceBound, brockschmidt2016AnalyzingRuntimeSize, Flores-MontoyaH14, giesl2022ImprovingAutomaticComplexity, hoffmann2017AutomaticResourceBound, lopez18IntervalBasedResource, carbonneaux2015CompositionalCertifiedResource, albert2012CostAnalysisObjectoriented, lommen2024ControlFlowRefinementProbabilistic, pham2024RobustResourceBounds, hoffmann2012MultivariateAmortizedResource, albert2013InferenceResourceUsage, handley2019LiquidateYourAssets], but most of them are essentially limited to non-recursive programs. There are also several techniques for complexity analysis of term rewrite systems (TRSs) [baader_nipkow_1998] which can handle arbitrary recursion. However, TRSs do not support built-in data types like integers. Thus, the goal of this paper is to automatically analyze the complexity of programs with built-in integers and arbitrary (possibly non-tail) recursion.
@@ 2.1 Syntax and Semantics of Recursive Integer Transition Systems
We also use $\geq$”, “==”, $\neq$”, and $\neg$”, since they can be simulated by atoms and constraints (e.g., $p_{1}\geq p_{2}$ is equivalent to $p_{2}<p_{1}+1$ for integers). Disjunctions $\lor$” are modeled by several transitions with the same start and target location.
@@
  • $\mathcal{T}$ is a finite set of transitions: A transition is a 4-tuple $(\ell,\varphi,\eta,\ell^{\prime})$ with start location $\ell\in\mathcal{L}\setminus\Omega$, target location $\ell^{\prime}\in\mathcal{L}\setminus\left\{\ell_{0}\right\}$, guard $\varphi\in\mathcal{C}$, and update function $\eta:\mathcal{V}\to\mathbb{Z}[\mathcal{V}\cup\mathcal{F}]$.
@@ 2.2 Runtime and Size Bounds for $\rho$-ITSs
The set of bounds $\mathcal{B}$ is the smallest set with $\overline{\mathbb{N}}\subseteq\mathcal{B}$, $\mathcal{V}\subseteq\mathcal{B}$, and $\{b_{1}+b_{2},\max(b_{1},b_{2}),b_{1}\cdot b_{2},p^{b_{1}},\log_{k}(b_{1})\}\subseteq\mathcal{B}$ for all $b_{1},b_{2}\in\mathcal{B}$, all polynomials $p\in\mathbb{N}[\mathcal{V}]$, and all $k\in\mathbb{R}_{>1}$. [^2] More precisely, instead of $\log_{k}(b_{1})$ we use $\lceil\log_{k}(\max\left\{1,b_{1}\right\})\rceil$ to ensure that bounds are well defined, weakly monotonically increasing, and evaluate to natural numbers.
@@
Our approach performs a modular analysis, i.e., parts of the program are analyzed as standalone programs and the results are then lifted to contribute to the overall analysis. So to compute a runtime bound for a transition $t$, our approach considers all transitions and function calls $\vartheta\in\mathcal{T}\cup\mathcal{F}$ that can occur directly before $t$ in evaluations, and it needs size bounds ${\mathcal{SB}}(\vartheta,v)$ to over-approximate the absolute values that the variables $v\in\mathcal{V}$ may have after these “previous” transitions and function calls $\vartheta$. We call $\mathcal{RV}=(\mathcal{T}\cup\mathcal{F})\times\mathcal{V}$ the set of result variables. Note that in contrast to runtime bounds (and to our earlier papers), we now also have to capture the effect of function calls $\mathcal{F}$ via size bounds.
@@ 3.2 Lifting Local Runtime Bounds to Global Runtime Bounds
To lift local to global runtime bounds, we consider those transitions and function calls which start an evaluation of the subprogram $\mathcal{T}^{\prime}$. Remember that $\operatorname{fun}(\mathcal{L}_{\mathcal{T}^{\prime}})$ denotes the set of all function calls $\ell(\,\underline{\hskip 4.26773pt}\,)\in\mathcal{F}$ with $\ell\in\mathcal{L}_{\mathcal{T}^{\prime}}$, where $\mathcal{L}_{\mathcal{T}^{\prime}}$ are the locations in the subprogram $\mathcal{T}^{\prime}$.
@@ 4 Modular Computation of Size Bounds
For every result variable $\langle t,v\rangle\in\mathcal{T}\times\mathcal{V}$ with $t=(\,\underline{\hskip 4.26773pt},\underline{\hskip 4.26773pt},\eta,\underline{\hskip 4.26773pt}\,)$, in practice we essentially use ${\mathcal{SB}_{\text{loc}}}(t,v)=\lceil\!\!\lceil\eta(v)\rceil\!\!\rceil$, e.g., ${\mathcal{SB}_{\text{loc}}}(t_{1},y)=y+\rho_{1}$ and ${\mathcal{SB}_{\text{loc}}}(t_{1},x)=\lceil\!\!\lceil x-1\rceil\!\!\rceil=x+1$ for the program from Fig. 2. However, due to the guard $x>0$ of $t_{1}$, here we can obtain the more precise local size bound ${\mathcal{SB}_{\text{loc}}}(t_{1},x)=x$. Similarly, we essentially use ${\mathcal{SB}_{\text{loc}}}(\rho,v)=\lceil\!\!\lceil\zeta(v)\rceil\!\!\rceil$ for every result variable $\langle\rho,v\rangle\in\mathcal{F}\times\mathcal{V}$ with $\rho=\ell(\zeta)$. So for Fig. 2, we would obtain ${\mathcal{SB}_{\text{loc}}}(\rho_{2},a)=\lceil\!\!\lceil a-1\rceil\!\!\rceil=a+1$. However, due to the guard $a>0$ of $t_{5}$ whose update contains the function call $\rho_{2}$, we can again obtain the more precise bound ${\mathcal{SB}_{\text{loc}}}(\rho_{2},a)=a$.
@@ 0.A.1 Proof of Lemma 1
$\leq\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{0}-i,n_{2}-1)\right)+\mathcal{R}_{n_{0}}(0,n_{2}).$ (6)
@@
$\leq\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{1}-i,n_{2}-1)\right)+\mathcal{R}_{n_{0}}(0,n_{2}).$ (by performing these steps repeatedly)
@@
$\mathcal{R}_{n_{0}}(n_{1},n_{2})$ $\leq\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{1}-i,n_{2}-1)\right)+\mathcal{R}_{n_{0}}(0,n_{2})$ (by (6))
@@
Let $\mathbb{T}_{1},\ldots,\mathbb{T}_{m}$ be the maximal subtrees of $\mathbb{T}$ where all edges are labeled with transitions from $\mathcal{T}^{\prime}$ or function calls from $\operatorname{fun}(\mathcal{T}^{\prime})\cap\operatorname{fun}(\mathcal{L}_{\mathcal{T}^{\prime}})$. So, the subtrees $\mathbb{T}_{1},\ldots,\mathbb{T}_{m}$ are constructed by only using $\prec_{\mathcal{T}^{\prime}\cup\left\{\varepsilon\right\}}$-steps, i.e., if $\mathbb{T}_{i}$’s root node is labeled with $(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})$, then we have $(\{(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})\},\varnothing)\prec_{\mathcal{T}^{\prime}\cup\left\{\varepsilon\right\}}^{*}\mathbb{T}_{i}$. Let $\vartheta_{i}\in\mathcal{T}\cup\mathcal{F}$ be the transition or the function call that the edge to $(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})$ is labeled with in $\mathbb{T}$, i.e., $\vartheta_{i}$ starts the evaluation of the subprogram $\mathcal{T}^{\prime}$. Let $k_{i}$ be the number of edges labeled with $t$ in $\mathbb{T}_{i}$ and let $\mathbb{T}$ have $k$ edges labeled with $t$, i.e., $\lvert\mathbb{T}_{i}\rvert_{\left\{t\right\}}=k_{i}$ and $\lvert\mathbb{T}\rvert_{\left\{t\right\}}=k$. Then we have $\sum_{i=1}^{m}k_{i}=k$.
@@
So let us consider $\alpha=\langle\vartheta,x\rangle\in C$. Note that $\vartheta$ cannot be an initial transition as there are no transitions or function calls leading back to the initial location $\ell_{0}$ (i.e., then $C$ would not be a non-trivial SCC). Hence, there exists a predecessor $\tilde{\vartheta}$ of $\vartheta$ in the path, i.e., the path has the form
@@
$=\llbracket s_{\alpha}\rrbracket_{\lvert\tilde{\sigma}\rvert}\cdot\left(e_{\alpha}+\sum_{v\in V_{\alpha}}\lvert\tilde{\sigma}(v)\rvert+\sum\limits_{\begin{subarray}{c}v\,\in\,\operatorname{actV}({\mathcal{SB}_{\text{loc}}}(\alpha))\\ v\,\not\in\,\operatorname{actV}(s_{\alpha})\cup V_{\alpha}\end{subarray}}\hskip-8.5359pt\llbracket init_{\alpha}(v)\rrbracket_{\lvert\sigma_{0}\rvert}\right$.

### v2

@@ 1 Introduction
There exist numerous approaches to analyze complexity of programs automatically, e.g., [ben-amram2017MultiphaseLinearRankingFunctions, albert2019ResourceAnalysisDriven, sinn2017ComplexityResourceBound, brockschmidt2016AnalyzingRuntimeSize, Flores-MontoyaH14, giesl2022ImprovingAutomaticComplexity, hoffmann2017AutomaticResourceBound, lopez18IntervalBasedResource, carbonneaux2015CompositionalCertifiedResource, albert2012CostAnalysisObjectoriented, lommen2024ControlFlowRefinementProbabilistic, pham2024RobustResourceBounds, hoffmann2012MultivariateAmortizedResource, albert2013InferenceResourceUsage, handley2019LiquidateYourAssets], but most of them are essentially limited to non-recursive programs. There are also several techniques for complexity analysis of term rewrite systems (TRSs) [baader_nipkow_1998] which can handle arbitrary recursion. However, TRSs have the drawback that they do not support built-in data types like integers. Thus, the goal of this paper is to automatically analyze the complexity of programs with built-in integers and arbitrary (possibly non-tail) recursion.
@@ 2.1 Syntax and Semantics of Recursive Integer Transition Systems
We also use $\geq$”, “==”, $\neq$”, and negations $\neg$”, since they can be simulated by atoms and constraints (e.g., $p_{1}\geq p_{2}$ is equivalent to $p_{2}<p_{1}+1$ for integers). Disjunctions $\lor$” are modeled by several transitions with the same start and target location.
@@
  • $\mathcal{T}$ is a finite set of transitions: A transition is a 4-tuple $(\ell,\varphi,\eta,\ell^{\prime})$ with start location $\ell\in\mathcal{L}$, target location $\ell^{\prime}\in\mathcal{L}\setminus\left\{\ell_{0}\right\}$, guard $\varphi\in\mathcal{C}$, and update function $\eta:\mathcal{V}\to\mathbb{Z}[\mathcal{V}\cup\mathcal{F}]$.
@@ 2.2 Runtime and Size Bounds for $\rho$-ITSs
The set of bounds $\mathcal{B}$ is the smallest set with $\overline{\mathbb{N}}\subseteq\mathcal{B}$, $\mathcal{V}\subseteq\mathcal{B}$, and $\{b_{1}+b_{2},\max(b_{1},b_{2}),b_{1}\cdot b_{2},p^{b_{1}},\log_{k}(b_{1})\}\subseteq\mathcal{B}$ for all $b_{1},b_{2}\in\mathcal{B}$, all polynomials $p\in\mathbb{N}[\mathcal{V}]$, and all $k\in\mathbb{R}_{>1}$. [^2] More precisely, instead of $\log_{k}(b_{1})$ we use the function $\lceil\log_{k}(\max\left\{1,b_{1}\right\})\rceil$ to ensure that bounds are well defined, weakly monotonically increasing, and evaluate to natural numbers.
@@
Our approach performs a modular analysis, i.e., parts of the program are analyzed as standalone programs and the results are then lifted to contribute to the overall analysis. So to compute a runtime bound for a transition $t$, our approach considers all transitions and function calls $\vartheta\in\mathcal{T}\cup\mathcal{F}$ that can occur before $t$ in evaluations, and it needs size bounds ${\mathcal{SB}}(\vartheta,v)$ to over-approximate the absolute values that the variables $v\in\mathcal{V}$ may have after these “previous” transitions and function calls $\vartheta$. We call $\mathcal{RV}=(\mathcal{T}\cup\mathcal{F})\times\mathcal{V}$ the set of result variables. Note that in contrast to runtime bounds (and to our earlier papers), we now also have to capture the effect of function calls $\mathcal{F}$ via size bounds.
@@ 3.2 Lifting Local Runtime Bounds to Global Runtime Bounds
To lift local to global runtime bounds, we consider those transitions and function calls which start an evaluation of the subprogram $\mathcal{T}^{\prime}$. Remember that $\operatorname{fun}(\mathcal{L}_{\mathcal{T}^{\prime}})$ denotes the set of all function calls $\ell(\,\underline{\hskip 4.26773pt}\,)\in\mathcal{F}$ with $\ell\in\mathcal{L}_{\mathcal{T}^{\prime}}$, where $\mathcal{L}_{\mathcal{T}^{\prime}}$ are the start locations of the transitions in the subprogram $\mathcal{T}^{\prime}$.
@@ 4 Modular Computation of Size Bounds
For every result variable $\langle t,v\rangle\in\mathcal{T}\times\mathcal{V}$ with $t=(\,\underline{\hskip 4.26773pt},\underline{\hskip 4.26773pt},\eta,\underline{\hskip 4.26773pt}\,)$, in practice we essentially use ${\mathcal{SB}_{\text{loc}}}(t,v)=\lceil\!\!\lceil\eta(v)\rceil\!\!\rceil$, e.g., ${\mathcal{SB}_{\text{loc}}}(t_{1},y)=y+\rho_{1}$ and ${\mathcal{SB}_{\text{loc}}}(t_{1},x)=\lceil\!\!\lceil x-1\rceil\!\!\rceil=x+1$ for the program from Fig. 2. However, due to the guard $x>0$ of the transition $t_{1}$, here we can obtain the more precise local size bound ${\mathcal{SB}_{\text{loc}}}(t_{1},x)=x$. Similarly, we essentially use ${\mathcal{SB}_{\text{loc}}}(\rho,v)=\lceil\!\!\lceil\zeta(v)\rceil\!\!\rceil$ for every result variable $\langle\rho,v\rangle\in\mathcal{F}\times\mathcal{V}$ with $\rho=\ell(\zeta)$. So for Fig. 2, we would obtain ${\mathcal{SB}_{\text{loc}}}(\rho_{2},a)=\lceil\!\!\lceil a-1\rceil\!\!\rceil=a+1$. However, due to the guard $a>0$ of the transition $t_{5}$ whose update contains the function call $\rho_{2}$, we can again obtain the more precise bound ${\mathcal{SB}_{\text{loc}}}(\rho_{2},a)=a$.
@@ 0.A.1 Proof of Lemma 1
$\leq\mathcal{R}_{n_{0}}(0,n_{2})+\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{0}-i,n_{2}-1)\right).$ (6)
@@
$\leq\mathcal{R}_{n_{0}}(0,n_{2})+\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{1}-i,n_{2}-1)\right).$ (by performing these steps repeatedly)
@@
$\mathcal{R}_{n_{0}}(n_{1},n_{2})$ $\leq\mathcal{R}_{n_{0}}(0,n_{2})+\sum_{i=0}^{n_{1}-1}\left(1+n_{0}+f(n_{0},n_{1}-i,n_{2}-1)\right)$ (by (6))
@@
Let $\mathbb{T}_{1},\ldots,\mathbb{T}_{m}$ be the maximal subtrees of $\mathbb{T}$ where all edges are labeled with transitions from $\mathcal{T}^{\prime}$ or function calls from $\operatorname{fun}(\mathcal{T}^{\prime})\cap\operatorname{fun}(\mathcal{L}_{\mathcal{T}^{\prime}})$. So, the subtrees $\mathbb{T}_{1},\ldots,\mathbb{T}_{m}$ are constructed by only using $\prec_{\mathcal{T}^{\prime}\cup\left\{\varepsilon\right\}}$-steps, i.e., if $\mathbb{T}_{i}$’s root node is labeled with $(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})$, then we have $(\{(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})\},\varnothing)\prec_{\mathcal{T}^{\prime}\cup\left\{\varepsilon\right\}}^{*}\mathbb{T}_{i}$. Let $\vartheta_{i}\in\mathcal{T}\cup\mathcal{F}$ be the transition or the function call in the label of the edge to $(\widetilde{\ell_{i}},{\tilde{\sigma}}_{i})$ in $\mathbb{T}$, i.e., $\vartheta_{i}$ starts the evaluation of the subprogram $\mathcal{T}^{\prime}$. Let $k_{i}$ be the number of edges labeled with $t$ in $\mathbb{T}_{i}$ and let $\mathbb{T}$ have $k$ edges labeled with $t$, i.e., $\lvert\mathbb{T}_{i}\rvert_{\left\{t\right\}}=k_{i}$ and $\lvert\mathbb{T}\rvert_{\left\{t\right\}}=k$. Then we have $\sum_{i=1}^{m}k_{i}=k$.
@@
So let us consider $\alpha=\langle\vartheta,x\rangle\in C$. Note that $\vartheta$ cannot be an initial transition as there are no transitions or function calls leading back to the initial location $\ell_{0}$ (i.e., then $C$ would not be a non-trivial SCC). Hence, there exists a predecessor
$\tilde{\vartheta}$ of $\vartheta$ in the path, i.e., the path has the form
@@
$=\llbracket s_{\alpha}\rrbracket_{\lvert\tilde{\sigma}\rvert}\cdot\left(e_{\alpha}+\sum_{v\in V_{\alpha}}\lvert\tilde{\sigma}(v)\rvert+\sum\limits_{\begin{subarray}{c}v\,\in\,\operatorname{actV}({\mathcal{SB}_{\text{loc}}}(\alpha))\\ v\,\not\in\,\operatorname{actV}(s_{\alpha})\cup V_{\alpha}\end{subarray}}\hskip-19.91684pt\llbracket init_{\alpha}(v)\rrbracket_{\lvert\sigma_{0}\rvert}\right$.